A particle has a kinetic energy equal three-quarters its rest mass energy. What is the speed of this particle? (Answer in terms of c)
Answer #1

KE = (γ - 1)*mo*c^2 = 0.75*mo*c^2

γ = 1.75 = [1 – (u/c)^2]-1/2

u = c*sqrt[1 – (1/γ)^2] = c*sqrt(0.6735) = 0.821*c